OpenStudy (anonymous):

Integrate : sinx - sin2x

OpenStudy (anonymous):

-cos(x)+cos(2x)/2

OpenStudy (anonymous):

rewrite sin2x as 2sinxcosx

OpenStudy (anonymous):

yes, i did that i thought i would have to use substitution

OpenStudy (anonymous):

yeah u-sub should take care of it

OpenStudy (anonymous):

or better yet, just remember it because math teachers seem to love this one. \[\int\sin(2x)dx=-\frac{1}{2}\cos(2x)\] \[\int \ln(x)dx=x\ln(x)-x\] save you fifteen minutes an exam really

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

ill try both ways....la grange, would i take out a sinx once i rewrite the equation?

OpenStudy (anonymous):

you could or you could simply write two integrals

OpenStudy (anonymous):